# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v2

@@ IX Discussion
While text generation and comprehension models are also typically large, manual inspection revealed that the ONNX Optimizer applies only a limited set of optimizations to them. Most passes are skipped, as many of the layer patterns targeted by the optimizer (e.g., convolutions)—common in classification and object detection models—are not typically present in transformer-based architectures used for language tasks.
(2) Optimization can reduce the accuracy of object detection models.
While some object detection models remained relatively stable after optimization, others showed notable performance degradation. For example, the YOLOv3 model maintained high detection accuracy with no misclassifications, although it exhibited minor shifts in bounding box positions. YOLOv2 also showed stable performance, with slightly lower Intersection over Union (IoU) due to its use of multiple bounding boxes per object, but still achieved strong detection metrics overall.
In contrast, the SSD model experienced a significant decline in F1 score, precision, and recall, leading to higher misclassification rates despite retaining a high IoU. Across all models, we observed substantial drops in mean Average Precision (mAP), with reductions of up to 14.31%, suggesting that optimization may cause certain object classes to be missed. This raises concerns about the reliability of optimized models in object detection and segmentation tasks.
(3) Optimizations can affect label accuracy in classification models. We identified two primary causes for differences in label accuracy following optimization. (a) The first relates to the implicit and improper upgrade of the ONNX format version in models using older opset versions. A significant number of classification models were built on these legacy opsets, making them particularly susceptible to this issue. (b) The second involves the generation of problematic graph nodes, particularly in relation to the fuse_bn_into_conv pass. This pass, which fuses batch normalization into convolutional layers, has a strong impact on classification models and can lead to unintended changes in model behavior after optimization.
(4) Outputs from Text comprehension and generation models were not affected by optimization. From the $8$ model instances under test related to text comprehension and generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model.
@@
X Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected $9.2$% of model instances that crashed when using the primary optimizer strategies, while $30$% of classification models and $16.6$% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified $15$ previously unreported bugs related to crashes and accuracy degradation across $9$ optimizer passes, which we reported to the ONNX Optimizer developers.
XI Availability
For more information about OODTE, please contact the authors. We intend to release the tool’s source code in the near future.

### v3

@@ 9. Discussion
(1) Optimizations can affect label accuracy in classification models. We identified two primary causes for label accuracy differences after optimization: (a) Legacy opset upgrade: Some models used older ONNX opsets, and the optimizer implicitly and improperly upgraded their format versions. Classification models built on these legacy opsets were particularly affected. (b) Problematic graph node generation: The fuse_bn_into_conv pass, which fuses batch normalization into convolutional layers, can unintentionally alter model behavior and particularly affects classification models that rely on convolutions.
Although text generation and comprehension models are also large, manual inspection showed that the optimizer applies only a few passes to them. Most optimizations are skipped because the targeted layer patterns (e.g., convolutions) are uncommon in transformer-based architectures used for language tasks.
@@
(3) Optimization can affect bounding boxes of object detection models. Regarding accuracy, object detection models remained largely stable. We observed only a negligible decrease in the accuracy of SSD on top-10 accuracy predictions (0.040.04%), while the other models showed no difference between the original and optimized versions. However, the optimized models occasionally misaligned bounding boxes, leading to drops in IoU value of up to 15.1915.19%.
(4) Outputs from text comprehension and generation models were not affected by optimization. Among the 8 text comprehension and generation models tested, we observed only one crash (in GPT-2) and a small number of samples (8) with BLEU-score differences between the original and optimized models for the top-3 token selections.
@@
10. Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected 9.29.2% of model instances that crashed when using the primary optimizer strategies, while 3030% of classification models and 16.616.6% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.